# Enhancing Consistency and Mitigating Bias: A Data Replay Approach for Incremental Learning

### v1

@@ VI Conclusion
In this paper, we propose CCIL framework to solve the catastrophic forgetting problem in data-free incremental learning. To address the data inconsistency problem in the literature, we first propose quantitative measures of data consistency. With them, we further analyze existing losses, which gives several insights and inspires the development of a novel loss term. Specifically, by aligning the statistical parameters in the feature space, we narrow the gap between synthetic and real data and ameliorate the inversion stage. This approach proves to be easy to implement. Furthermore, we identify a phenomenon where the norms of old class weights decrease as learning progresses. We analyze the underlying reasons in the background of DFCIL and propose a simple and effective regularization term to reduce the weight bias. Experiments on different datasets show our method can surpass previous works and achieve SOTA performance. Also importantly, we believe the comprehensive analysis of multiple aspects of data-free data replay methods in our study will contribute to the ongoing efforts in developing more effective techniques.

### v2

@@ 7 Conclusion
In this paper, we propose CwD framework to solve the catastrophic forgetting problem in data-free incremental learning. To address the data inconsistency problem in the literature, we first propose the quantitative measure of data consistency, which further inspires the development of a novel loss term. Specifically, by aligning the statistical parameters in the feature space, we narrow the gap between synthetic and real data and ameliorate the inversion stage. This approach proves to be easy to implement. Furthermore, we identify a phenomenon where the norms of old class weights decrease as learning progresses. We analyze the underlying reasons in the background of DFCIL and propose a simple and effective regularization term to reduce the weight bias. Experiments on different datasets show our method can surpass previous works and achieve SOTA performance. Also importantly, we believe the comprehensive analysis of multiple aspects of data-free data replay methods in our study will contribute to the ongoing efforts to develop more effective techniques.
Acknowledgement
The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).